# Bias Similarity Across Large Language Models

### v3

@@ 7 Limitation
First, our evaluations are constrained by the available datasets, which cover only a subset of demographic dimensions—primarily gender, nationality, ethnicity, and religion—and are entirely in English. While we use all dimensions present in BBQ and UnQover, their overlap is partial and excludes axes like disability or intersectional biases. In addition, these benchmarks may not capture more subtle forms of bias, such as microaggressions or context-dependent harms that emerge over longer conversations. Furthermore, limiting the analysis to English overlooks how bias manifests in multilingual or code-switched contexts. Broader demographic coverage and cross-lingual evaluations are essential to assess global model fairness.
@@
Second, although we expand beyond multiple-choice QA using open-ended prompts from StereoSet (Appendix G), this evaluation remains limited in scope. Models often fail to generate valid completions, and even successful outputs vary greatly in structure. Our sentiment-based framing bias analysis captures only one aspect (i.e., sentiment) and does not account for deeper representational harm, refusal strategies, or evasive completions. Future work should expand bias evaluations to more interactive settings, such as multi-turn dialogue or retrieval-augmented tasks, where contextual harms may emerge more clearly.
@@
Finally, while we report a range of evaluation metrics—accuracy, bias scores, output histograms, flip statistics, cosine distance, Jensen-Shannon divergence, and CKA—across 24 LLMs and analyze the similarity between base and instruction-tuned models, we do not examine how these similarity patterns would change under alternative debiasing strategies. Techniques such as data augmentation, adversarial training, and representation-level debiasing may alter model behavior and internal representations in distinct ways, potentially leading to different similarity dynamics. However, our study focuses on naturally occurring behaviors in widely used foundation and instruction-tuned models, leaving the impact of targeted debiasing interventions as a valuable direction for future work.
8 Societal Impact and Ethical Consideration
Positive Impacts.
The improved bias assessment offers a strong foundation for advancing fairness in LLMs. By evaluating models across multiple contexts (ambiguous, disambiguated, and forced-choice), the framework captures deeper behavioral tendencies and quantifies the impact of mitigation efforts. It reveals that certain biases persist across model families and tuning strategies, pointing to structural patterns rooted in pretraining data or architecture. These insights support mitigation strategies beyond abstention—such as dataset balancing or representation-level debiasing—that meaningfully reduce directional bias. The framework also uncovers over-abstention, where models default to “unknown” even when clarity is possible. Recognizing this enables the design of models that are not only safer but also more contextually aware and practically useful. The finding that open-source models can match or exceed proprietary ones in fairness further promotes accessibility and transparency. Finally, by linking behavioral patterns with internal representations (e.g., via CKA), the framework supports multi-layered, behaviorally grounded auditing tools and provides a reproducible map for comparing models across scales and families.
Negative Impacts and Risks.
The findings carry significant societal implications. Persistent directional biases in forced-choice settings underscore the risk of LLMs subtly reinforcing harmful stereotypes. Meanwhile, the tendency of proprietary models to abstain, particularly in ambiguous contexts, can have uneven effects across applications, potentially erasing diversity or normalizing biased assumptions. In high-stakes domains such as healthcare or law, consistently responding with “unknown” to questions involving marginalized groups—despite clear contextual cues—may perpetuate informational inequity by withholding critical knowledge. These behaviors are also vulnerable to dual-use exploitation: malicious actors could craft prompts to bypass abstention filters or amplify biased outputs for misinformation, propaganda, or targeted persuasion.
Failure Modes.
Bias mitigation strategies that rely solely on abstention or instruction tuning may offer a false sense of safety. Our results show that models with high representational similarity can still diverge in behavior, producing biased outputs under pressure. Such failure modes are especially harmful for marginalized groups who may be poorly represented in training data or benchmarks. Without multi-metric, context-aware audits, developers risk deploying models that appear fair but behave unfairly in real-world use.
Appendix A Bias Definition and Illustrative Examples
In Table 2, we include the definition of bias across the four dimensions—gender, race, nationality, religion—and their illustrative examples.
@@ Appendix A Bias Definition and Illustrative Examples
Gender Associating certain behaviors, traits, or professions with specific genders (e.g., predicting males for leadership roles).
Race Linking certain races to particular roles or attributes (e.g., associating criminality with specific racial group).
Nationality Stereotyping individuals based on national origin (e.g., associating wealth with certain nations).
Religion Making assumptions based on religious affiliation (e.g., attributing violent tendencies to a particular faith).
Appendix B Detailed Analysis of Flip Behavior and Bias Scores
Abstention Trends and Effective Debiasing
Instruction tuning often increases “Unknown” predictions via S→U and A→U flips—desirable behavior in ambiguous prompts. The most effective debiasing cases are Gemma 2 9B-It, Gemma 2 27B-It, and Gemma 3 12B-It, each achieving over 50% abstention rates overall. For instance, Gemma 2 9B-It records a 73.1% UNK flip rate in gender and 60.5% in religion, with minimal retention (< 5%) or directional reversals. These models exhibit near-zero s_AMB, validating that abstention aligns with fairness-promoting moderation of directional bias.
Low Abstention and Bias Retention
In contrast, LLaMA 2 7B and Gemma 7B display low abstention (11.2-27.8%) and high retention of biased predictions (Ret(S) $>$ 60%). Their bias scores remain positive in both contexts, especially in nationality and religion. This suggests they often maintain or redistribute bias rather than neutralize it.
Unintended Reversals and Tuning Instability
Although some tuned models demonstrate increased abstention, they often introduce substantial directional flips. For instance, LLaMA 3 8B-Chat flips 118 anti-stereotypical (A→S) responses and 49 in the reverse (S→A) for gender, retaining 21% of biased outputs. Similarly, Gemma 3 4B-It introduces 386 A→S flips in gender while retaining $>$ 50% of stereotypes across dimensions, leading to increased s_DIS scores (e.g., gender: 2.69 → 8.62). These cases highlight how abstention gains can coexist with backsliding on fairness when directional reversals persist.
Scaling and Consistency
Model scale does not uniformly predict fairness gains. Gemma 3 12B-It exhibits more consistent improvement than its 27B variant, which shows higher A→S flips and stereotype retention despite similar abstention. Likewise, LLaMA 3 70B-Chat underperforms its 8B counterpart in flip rate (e.g., 14.2% vs. 34.9% in nationality), despite showing comparable s_DIS. It confirms that scaling alone does not determine debiasing success.
Summary and Insights
The bias scores and flip rates underscore the following key points:
Model Pair Dimension Total A→S S→A A→U S→U Ret(A) Ret(S) UNK Flip
LLaMA 2 7B → LLaMA 2 7B-Chat Ethnicity 3440 76 102 139 122 85.2 85.1 7.6
LLaMA 2 7B → LLaMA 2 7B-Chat Gender 2836 369 369 164 153 54.2 55.7 11.2
LLaMA 2 7B → LLaMA 2 7B-Chat Nationality 1540 0 0 70 54 89.3 92.0 8.1
LLaMA 2 7B → LLaMA 2 7B-Chat Religion 600 84 82 31 27 54.9 58.1 9.7
LLaMA 3 8B → LLaMA 3 8B-Chat Ethnicity 3440 27 12 727 843 36.9 28.2 45.6
LLaMA 3 8B → LLaMA 3 8B-Chat Gender 2836 118 49 462 557 21.0 36.5 35.9
LLaMA 3 8B → LLaMA 3 8B-Chat Nationality 1540 0 0 215 323 61.5 40.5 34.9
LLaMA 3 8B → LLaMA 3 8B-Chat Religion 600 31 15 103 132 23.9 34.7 39.2
LLaMA 3 70B → LLaMA 3 70B-Chat Ethnicity 3440 0 0 340 376 38.4 35.7 20.8
LLaMA 3 70B → LLaMA 3 70B-Chat Gender 2836 38 20 133 122 34.5 55.2 9.0
LLaMA 3 70B → LLaMA 3 70B-Chat Nationality 1540 0 0 99 119 65.6 52.0 14.2
LLaMA 3 70B → LLaMA 3 70B-Chat Religion 600 10 11 29 58 30.4 50.0 14.5
Gemma 7B → Gemma 7B-It Ethnicity 3440 53 41 375 418 64.2 67.2 23.1
Gemma 7B → Gemma 7B-It Gender 2836 261 138 269 245 42.8 63.8 18.1
Gemma 7B → Gemma 7B-It Nationality 1540 0 0 194 214 67.4 67.7 26.5
Gemma 7B → Gemma 7B-It Religion 600 62 28 75 92 36.3 49.4 27.8
Gemma 2 9B → Gemma 2 9B-It Ethnicity 3440 0 0 1021 1126 4.3 4.4 62.4
Gemma 2 9B → Gemma 2 9B-It Gender 2836 1 0 954 1120 1.1 0.4 73.1
Gemma 2 9B → Gemma 2 9B-It Nationality 1540 0 0 396 514 20.3 6.2 59.1
Gemma 2 9B → Gemma 2 9B-It Religion 600 4 3 150 213 0.6 13.3 60.5
Gemma 2 27B → Gemma 2 27B-It Ethnicity 3440 0 0 819 928 8.8 9.0 50.8
Gemma 2 27B → Gemma 2 27B-It Gender 2836 1 0 709 937 0.0 0.4 58.0
Gemma 2 27B → Gemma 2 27B-It Nationality 1540 0 0 217 426 34.4 5.1 41.8
Gemma 2 27B → Gemma 2 27B-It Religion 600 8 4 114 122 4.7 22.2 39.3
Gemma 3 4B → Gemma 3 4B-It Ethnicity 3440 46 41 660 570 58.1 64.2 35.8
Gemma 3 4B → Gemma 3 4B-It Gender 2836 386 171 484 521 33.1 53.6 35.4
Gemma 3 4B → Gemma 3 4B-It Nationality 1540 0 0 203 231 70.9 68.7 28.2
Gemma 3 4B → Gemma 3 4B-It Religion 600 81 38 104 160 31.7 36.9 44.0
Gemma 3 12B → Gemma 3 12B-It Ethnicity 3440 1 2 927 1058 15.0 12.3 57.7
Gemma 3 12B → Gemma 3 12B-It Gender 2836 55 19 683 829 5.4 14.1 53.3
Gemma 3 12B → Gemma 3 12B-It Nationality 1540 0 0 225 408 41.6 16.0 41.1
Gemma 3 12B → Gemma 3 12B-It Religion 600 17 4 107 150 12.1 27.7 42.8
Gemma 3 27B → Gemma 3 27B-It Ethnicity 3440 1 3 793 852 5.9 6.5 47.8
Gemma 3 27B → Gemma 3 27B-It Gender 2836 7 3 548 653 1.2 6.3 42.3
Gemma 3 27B → Gemma 3 27B-It Nationality 1540 0 0 366 449 25.3 8.2 52.9
Gemma 3 27B → Gemma 3 27B-It Religion 600 9 2 122 154 0.0 19.6 46.0
@@ Appendix B Detailed Analysis of Flip Behavior and Bias Scores
Table 4: Bias scores for ambiguous and disambiguated questions across four dimensions. Scores near 0 indicate neutrality; positive and negative values reflect stereo- and anti-stereotypical bias. Large drops between s_DIS and s_AMB suggest correct abstention in ambiguous settings but directional bias when models are forced to choose.
@@
Gender Nationality Ethnicity Religion Gender Nationality Ethnicity Religion
@@
Appendix C Response Histograms
Figure 5: BBQ Response Distribution Histograms. Each figure shows responses distribution to ambiguous prompts in BBQ, broken down by bias dimensions. While “Unknown” is often the dominant response, it is less prevalent in certain underrepresented dimensions, such as age, ses, or disability, revealing variation in abstention behavior.
@@ Appendix C Response Histograms
Appendix D Cosine Similarity Matrices (Detailed)
Low and Consistent Distances in BBQ
Figure 7 shows that cosine distances in the ambiguous BBQ setting are generally low and consistent across dimensions, indicating modest tuning effects on directional output behavior. The standout outlier is Gemma 3 4B vs. 4B-It (0.58), consistent with its large abstention shift observed in Figure 5. Aside from this, distances remain tightly clustered, even across families, such as LLaMA 3 and Gemma 3.
Greater Dimensional Variability in UnQover
UnQover exhibits more variability in model behavior, particularly across dimensions. Ethnicity and religion show relatively stable distance patterns, while gender and nationality produce more dispersed cosine distances, indicating greater divergence in model preferences.
Gemma 3 27B-It and Gemini 1.5/2.0 frequently appear as outliers, exhibiting high dissimilarity from all other models—and occasionally from one another. They align in some dimensions (e.g., ethnicity, religion) but diverge in others (e.g., gender, nationality). Gemma 2 9B-It also behaves inconsistently, sometimes clustering with tuned or proprietary models, sometimes not.
Histograms in Figure 6 clarify this pattern. The outlier models produce high counts of “Unknown” across all dimensions but differ in how they distribute remaining responses. Religion and ethnicity are relatively balanced; nationality and gender exhibit skew, which corresponds to the increased cosine distance variability.
This explains the observed variability in UnQover cosine distances: tuning effects and cautious responses manifest differently across dimensions. It reinforces our core claim that fairness behavior is context- and dimension-dependent, and that no single metric or prompt format fully captures how models handle bias.
@@ Appendix D Cosine Similarity Matrices (Detailed)
Appendix E JS Divergence Across Models and Dimensions

### v4

@@ Appendix A Limitation
First, our evaluations are constrained by the available datasets, which cover only a subset of demographic dimensions (e.g., primarily gender, nationality, ethnicity, and religion) and are entirely in English. While we use all dimensions present in BBQ and UnQover, their overlap is partial and excludes axes like disability or intersectional biases. These benchmarks also may not capture subtler forms of bias, such as microaggressions or context-dependent harms that emerge over longer interactions. In addition, limiting the analysis to English overlooks how bias manifests in multilingual or code-switched contexts. Broader demographic coverage and cross-lingual evaluations are essential for assessing global model fairness.
@@
Second, although we expand beyond multiple-choice QA using open-ended prompts from StereoSet (Appendix H), this evaluation remains limited in scope. Models often fail to generate valid completions, and even successful outputs vary greatly in structure. Our sentiment-based framing bias analysis captures only one aspect (sentiment polarity) and does not account for deeper representational harms, refusal strategies, or evasive completions. Future work should extend bias evaluation to more interactive settings, such as multi-turn dialogue or retrieval-augmented tasks, where contextual harms may emerge more clearly.
@@
Third, while we report a range of evaluation metrics (accuracy, bias scores, output histograms, flip statistics, cosine distance, JSD, and CKA) across 30 LLMs and analyze similarity under diverse conditions (base vs. tuned, release versions, model sizes, open vs. proprietary, and across families), we do not examine how these patterns would change under targeted debiasing strategies. Approaches such as data augmentation, adversarial training, or representation-level debiasing may alter model behavior and internal representations in distinct ways, potentially leading to different similarity dynamics. Our study instead focuses on naturally occurring behaviors in widely used models, leaving the effects of deliberate debiasing interventions as a valuable direction for future work.
Finally, our analysis is constrained by practical and methodological factors. Inference cost limited full coverage across datasets (e.g., GPT-5-mini was excluded on UnQover), and API-only models prevented deeper representation-level comparisons. Moreover, cross-vendor comparisons should be interpreted as descriptive ecosystem mapping rather than causal attribution, since architectures, data, and tuning pipelines differ in uncontrolled ways. These constraints highlight the need for complementary studies with broader resources and controlled settings.
@@
Gender Associating certain behaviors, traits, or professions with specific genders
(e.g., predicting males for leadership roles).
Race Linking certain races to particular roles or attributes
(e.g., associating criminality with a specific racial group).
Nationality Stereotyping individuals based on national origin
(e.g., associating wealth with certain nations).
Religion Making assumptions based on religious affiliation
(e.g., attributing violent tendencies to a particular faith).
Appendix B Societal Impact and Ethical Consideration
Positive Impacts. The improved bias assessment offers a strong foundation for advancing fairness in LLMs. By evaluating models across multiple contexts (ambiguous, disambiguated, and forced-choice), the framework captures deeper behavioral tendencies and quantifies the impact of mitigation efforts. It reveals that certain biases persist across model families and tuning strategies, pointing to structural patterns rooted in pretraining data or architecture. These insights support mitigation strategies beyond abstention—such as dataset balancing or representation-level debiasing—that meaningfully reduce directional bias. The framework also uncovers over-abstention, where models default to “unknown” even when clarity is possible. Recognizing this enables the design of models that are not only safer but also more contextually aware and practically useful. The finding that open-source models can match or exceed proprietary ones in fairness further promotes accessibility and transparency. Finally, by linking behavioral patterns with internal representations (e.g., via CKA), the framework supports multi-layered, behaviorally grounded auditing tools and provides a reproducible map for comparing models across scales and families.
Negative Impacts and Risks. The findings carry significant societal implications. Persistent directional biases in forced-choice settings underscore the risk of LLMs subtly reinforcing harmful stereotypes. Meanwhile, the tendency of proprietary models to abstain, particularly in ambiguous contexts, can have uneven effects across applications, potentially erasing diversity or normalizing biased assumptions. In high-stakes domains such as healthcare or law, consistently responding with “unknown” to questions involving marginalized groups—despite clear contextual cues—may perpetuate informational inequity by withholding critical knowledge. These behaviors are also vulnerable to dual-use exploitation: malicious actors could craft prompts to bypass abstention filters or amplify biased outputs for misinformation, propaganda, or targeted persuasion.
Failure Modes. Bias mitigation strategies that rely solely on abstention or instruction tuning may offer a false sense of safety. Our results show that models with high representational similarity can still diverge in behavior, producing biased outputs under pressure. Such failure modes are especially harmful for marginalized groups who may be poorly represented in training data or benchmarks. Without multi-metric, context-aware audits, developers risk deploying models that appear fair but behave unfairly in real-world use.
@@ Appendix B Societal Impact and Ethical Consideration
Table 5: Bias scores for ambiguous and disambiguated questions across four dimensions. Scores near 0 indicate neutrality; positive and negative values reflect stereo- and anti-stereotypical bias. Large drops between s_DIS and s_AMB suggest correct abstention in ambiguous settings but directional bias when models are forced to choose. Gen, Nat, Eth, and Rel refer to Gender, Nationality, Ethnicity, and Religion, respectively.
@@
Gen Nat Eth Rel Gen Nat Eth Rel
@@
LLaMA 3.1 8B 27.16 12.71 19.79 12.68 45.22 22.84 34.47 22.96
LLaMA 3.1 8B-Chat 2.31 2.18 1.04 0.00 8.45 6.76 3.91 0.00
LLaMA 3.1 70B 2.89 -0.76 -0.81 0.36 8.17 -2.01 -3.31 0.39
LLaMA 3.1 70B-Chat -0.34 0.02 -0.35 0.08 -2.82 0.32 -3.65 0.39
LLaMA 3.2 3B 25.30 6.76 9.99 5.77 36.30 10.47 14.90 9.03
LLaMA 3.2 3B-Chat 19.91 9.09 13.91 17.93 33.62 20.00 30.36 39.89
@@
GPT-5 Mini -0.00 0.11 -0.03 0.75 -0.21 1.82 -2.33 5.12
Model Pair Dimension Total A→S S→A A→U S→U Ret(A) Ret(S) UNK Flip
LLaMA 2 7B → Chat Ethnicity 3440 76 102 139 122 85.2 85.1 7.6
LLaMA 2 7B → Chat Gender 2836 369 369 164 153 54.2 55.7 11.2
LLaMA 2 7B → Chat Nationality 1540 0 0 70 54 89.3 92.0 8.1
LLaMA 2 7B → Chat Religion 600 84 82 31 27 54.9 58.1 9.7
LLaMA 3 8B → Chat Ethnicity 3440 27 12 727 843 36.9 28.2 45.6
LLaMA 3 8B → Chat Gender 2836 118 49 462 557 21.0 36.5 35.9
LLaMA 3 8B → Chat Nationality 1540 0 0 215 323 61.5 40.5 34.9
LLaMA 3 8B → Chat Religion 600 31 15 103 132 23.9 34.7 39.2
LLaMA 3 70B → Chat Ethnicity 3440 0 0 340 376 38.4 35.7 20.8
LLaMA 3 70B → Chat Gender 2836 38 20 133 122 34.5 55.2 9.0
LLaMA 3 70B → Chat Nationality 1540 0 0 99 119 65.6 52.0 14.2
LLaMA 3 70B → Chat Religion 600 10 11 29 58 30.4 50.0 14.5
LLaMA 3.1 8B → Chat Ethnicity 3440 11 11 779 929 34.1 34.4 49.7
LLaMA 3.1 8B → Chat Gender 2836 33 19 543 641 24.4 28.3 41.7
LLaMA 3.1 8B → Chat Nationality 1540 0 0 291 381 45.8 38.4 43.6
LLaMA 3.1 8B → Chat Religion 600 17 12 112 149 27.5 37.8 43.5
LLaMA 3.1 70B → Chat Ethnicity 3440 1 0 362 401 17.9 26.0 22.2
LLaMA 3.1 70B → Chat Gender 2836 12 22 178 247 32.1 27.5 15.0
LLaMA 3.1 70B → Chat Nationality 1540 0 0 284 297 29.9 19.5 37.7
LLaMA 3.1 70B → Chat Religion 600 7 3 67 133 7.5 37.0 33.3
LLaMA 3.2 3B → Chat Ethnicity 3440 21 13 874 912 44 42.9 51.9
LLaMA 3.2 3B → Chat Gender 2836 70 34 758 685 36.1 46.8 50.9
LLaMA 3.2 3B → Chat Nationality 1540 0 0 352 340 47.6 48.2 44.9
LLaMA 3.2 3B → Chat Religion 600 23 19 160 196 23.8 31.5 59.3
Gemma 7B → It Ethnicity 3440 53 41 375 418 64.2 67.2 23.1
Gemma 7B → It Gender 2836 261 138 269 245 42.8 63.8 18.1
Gemma 7B → It Nationality 1540 0 0 194 214 67.4 67.7 26.5
Gemma 7B → It Religion 600 62 28 75 92 36.3 49.4 27.8
Gemma 2 9B → It Ethnicity 3440 0 0 1021 1126 4.3 4.4 62.4
Gemma 2 9B → It Gender 2836 1 0 954 1120 1.1 0.4 73.1
Gemma 2 9B → It Nationality 1540 0 0 396 514 20.3 6.2 59.1
Gemma 2 9B → It Religion 600 4 3 150 213 0.6 13.3 60.5
Gemma 2 27B → It Ethnicity 3440 0 0 819 928 8.8 9.0 50.8
Gemma 2 27B → It Gender 2836 1 0 709 937 0.0 0.4 58.0
Gemma 2 27B → It Nationality 1540 0 0 217 426 34.4 5.1 41.8
Gemma 2 27B → It Religion 600 8 4 114 122 4.7 22.2 39.3
Gemma 3 4B → It Ethnicity 3440 46 41 660 570 58.1 64.2 35.8
Gemma 3 4B → It Gender 2836 386 171 484 521 33.1 53.6 35.4
Gemma 3 4B → It Nationality 1540 0 0 203 231 70.9 68.7 28.2
Gemma 3 4B → It Religion 600 81 38 104 160 31.7 36.9 44.0
Gemma 3 12B → It Ethnicity 3440 1 2 927 1058 15.0 12.3 57.7
Gemma 3 12B → It Gender 2836 55 19 683 829 5.4 14.1 53.3
Gemma 3 12B → It Nationality 1540 0 0 225 408 41.6 16.0 41.1
Gemma 3 12B → It Religion 600 17 4 107 150 12.1 27.7 42.8
Gemma 3 27B → It Ethnicity 3440 1 3 793 852 5.9 6.5 47.8
Gemma 3 27B → It Gender 2836 7 3 548 653 1.2 6.3 42.3
Gemma 3 27B → It Nationality 1540 0 0 366 449 25.3 8.2 52.9
Gemma 3 27B → It Religion 600 9 2 122 154 0.0 19.6 46.0
Appendix C Detailed Analysis of Flip Behavior and Bias Scores
Abstention Trends and Effective Debiasing. Instruction tuning often increases “Unknown” predictions via S→U and A→U flips, which is a desirable behavior in ambiguous prompts. The most effective debiasing cases are Gemma 2 9B-It, Gemma 2 27B-It, and Gemma 3 12B-It, each achieving over 50% abstention rates overall. For instance, Gemma 2 9B-It records a 73.1% UNK flip rate in gender and 60.5% in religion, with minimal retention ($5\%$) or directional reversals. These models exhibit near-zero s_AMB, validating that abstention aligns with fairness-promoting moderation of directional bias.
Low Abstention and Bias Retention. In contrast, LLaMA 2 7B and Gemma 7B display low abstention (11.2-27.8%) and high retention of biased predictions (Ret(S) >> 60%). Their bias scores remain positive in both contexts, especially in nationality and religion. This suggests they often maintain or redistribute bias rather than neutralize it.
Unintended Reversals and Tuning Instability. Although some tuned models demonstrate increased abstention, they often introduce substantial directional flips. For instance, LLaMA 3 8B-Chat flips 118 anti-stereotypical (A→S) responses and 49 in the reverse (S→A) for gender, retaining 21% of biased outputs. Similarly, Gemma 3 4B-It introduces 386 A→S flips in gender while retaining >> 50% of stereotypes across dimensions, leading to increased s_DIS scores (e.g., gender: 2.69 → 8.62). These cases highlight how abstention gains can coexist with backsliding on fairness when directional reversals persist.
Scaling and Consistency. Model scale does not uniformly predict fairness gains. Gemma 3 12B-It exhibits more consistent improvement than its 27B variant, which shows higher A→S flips and stereotype retention despite similar abstention. Likewise, LLaMA 3 70B-Chat underperforms its 8B counterpart in flip rate (e.g., 14.2% vs. 34.9% in nationality), despite showing comparable s_DIS. It confirms that scaling alone does not determine debiasing success.
Summary and Insights. The bias scores and flip rates underscore the following key points:
Appendix D Additional Response Histograms
@@
Figure 5: BBQ Response Distribution Histograms. Each figure shows the distribution of responses to ambiguous prompts in BBQ, broken down by bias dimensions. While “Unknown” is often the dominant response, it is less prevalent in certain underrepresented dimensions, such as age, ses, or disability, revealing variation in abstention behavior.
Figure 6: BBQ Response Distribution Histograms. Each figure shows the distribution of responses to ambiguous prompts in BBQ, broken down by bias dimensions. While “Unknown” is often the dominant response, it is less prevalent in certain underrepresented dimensions, such as age, ses, or disability, revealing variation in abstention behavior.
@@
Appendix E CKA Similarities Across Dimensions